When a consumer’s income increased from $54000 to $66000, they ate at a certain restaurant 36 times instead of 28. What is the income elasticity of demand in this case? Based on this calculated elasticity, what type of good is eating at this restaurant as far as this consumer is concerned?


1
Expert's answer
2020-07-13T13:36:49-0400

The income elasticity of demand in this case is:

"Ed = \frac{36 - 28} {66000-54000} × \frac{66000+54000} {36+28} = \frac{5} {4} = 1.25."

So, this is a normal good.